OpenStudy (anonymous):

What is the volume of a rectangular prism that is 120 centimeters by 2 meters by 1.5 meters in cubic centimeters? 3,600 cm^ 3,600,000 cm^ 36,000 cm^ 3.6 cm^

OpenStudy (anonymous):

120 times 2 times 1.5

Parth (parthkohli):

\( \color{Black}{\Rightarrow \color{Red}{length \times width \times height} }\) \( \color{Black}{\Rightarrow 120 \times 2 \times 1.5 }\)

OpenStudy (anonymous):

360

Parth (parthkohli):

Does telling the answer make it better? :/

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

sorry i thought he got it already

OpenStudy (anonymous):

what?